# Generative Topological Networks

### v1

@@ Abstract
Generative models have seen significant advancements in recent years, yet often remain challenging and costly to train and use. We introduce Generative Topological Networks (GTNs) – a new class of generative models that addresses these shortcomings. GTNs are trained deterministically using a simple supervised learning approach grounded in topology theory. GTNs are fast to train, and require only a single forward pass in a standard feedforward neural network to generate samples. We demonstrate the strengths of GTNs in several datasets, including MNIST, celebA and the Hands and Palm Images dataset. Finally, the theory behind GTNs offers insights into how to train generative models for improved performance. Code will be published upon acceptance.
@@ 1 Introduction
Figure 1: Samples generated from a standard normal distribution using a GTN trained on CelebA $64\times 64$ with a latent dimension of $100$.
Deep generative models such as Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), Variational Autoencoders (VAEs) (Kingma and Welling 2013), Energy-Based Models (EBMs) (LeCun et al. 2006; Ngiam et al. 2011), autoregressive models (Bengio and Bengio 2000; Larochelle and Murray 2011), normalizing flows (NFs) (Rezende and Mohamed 2015), and diffusion models (Sohl-Dickstein et al. 2015a), have demonstrated remarkable capabilities for generating samples based on training data distributions (Kang et al. 2023; Ho, Jain, and Abbeel 2020; Ramesh et al. 2021; Ho et al. 2019; Kingma and Dhariwal 2018; Kingma et al. 2016; Reed et al. 2017; Van Den Oord, Vinyals et al. 2017; Ramesh et al. 2021).
@@
However, diffusion models have several downsides. Most prominent is their inferior time-efficiency during training and sampling when compared to VAEs and GANs. This is because they employ a gradual and stochastic transition from a Gaussian distribution to the data distribution, performed over multiple time-steps known as the reverse diffusion process. The complexity of this process also makes it challenging to produce interpolations – desirable in the context of continuous data like audio and video. Additionally, diffusion-based methods often generate samples that fall outside the data distribution (e.g. the horizontally-flipped $3$s in (Sohl-Dickstein et al. 2015b) Figure App.1.). Although recent advancements have made significant steps towards enhancing diffusion models, the complex diffusion process is still a key challenging component (Manduchi et al. 2024).
@@
From a practical perspective, GTNs are extremely simple and efficient, requiring a single, vanilla, feedforward architecture trained using supervised learning. This means that GTNs avoid the complexities of introducing stochasticity during training which is present in VAE and diffusion; circumvent the intricacies of training more complex architectures such as those employed by diffusion and GANs; are not a likelihood-based approach, as are VAE, diffusion and NFs; and do not pose constraints on the structure of the neural network, as do NFs. The supervised manner in which GTNs are trained also forms a strong inductive bias that speeds up convergence. GTNs also operate on latent representations of the data, providing even more enhanced training and sampling efficiency. These advantages manifest in our experiments – realistic samples are obtained at early training epochs, and generating a single image takes around $4$ milliseconds on a standard PC and $0.003$ milliseconds on a T4 GPU.
@@
The remainder of the paper is structured as follows: In the Method section we first develop the theory behind GTNs for the $1$-dimensional case and use it to generate samples from the swiss-roll. We then extend this theory to higher dimensions and use it to generate samples from the multivariate uniform distribution. In the Experiments section we apply GTNs to real datasets, also including demonstrations of continuous interpolation. We follow with the Related Work section and conclude with the Discussion section.
@@ 2 Method
We first develop the method in the simple case of a $1$-dimensional space, and proceed to generalize to higher dimensions.
@@
In this section we will define a function $h$ that transforms one distribution into another (illustrated in Figure 2). We will also prove that $h$, under fairly general conditions, possesses certain properties that are desirable in the context of generative models by proving that it is a homeomorphism (Theorem 2.1). For example, we will see that $h$ is bijective, so that each sample $y$ is mapped precisely to one sample $x_{y}=h(y)$, generating different samples for different $y$, and guaranteeing that each sample $x$ has a sample $y$ that generates it. We will also discuss other useful consequences of $h$ being a homeomorphism. We begin by defining a homeomorphism in our context and proceed to define $h$ in Theorem 2.1.
Definition 2.1
(Homeomorphism for $\mathbb{R}^{n}$) Let $S,T$ be two subsets of $\mathbb{R}^{n}$. A function $h:S\to T$ is a homeomorphism if: (1) $h$ is continuous; (2) $h$ is bijective; (3) $h^{-1}$ is continuous. When such an $h$ exists, then $S$ and $T$ are called homeomorphic.
Theorem 2.1
Let $X$ and $Y$ be random variables that have continuous probability density functions (pdfs) $f_{X}$, $f_{Y}$ and supports $S_{X}$, $S_{Y}$ that are open intervals in $\mathbb{R}$. Denote the corresponding cumulative distribution functions (CDFs) as $F_{X}$ and $F_{Y}$. Define:
@@
Note that the requirement that $S_{X},S_{Y}$ are open intervals can be adapted to a union of open intervals – see Appendix A.3. The proof of Theorem 2.1 is in Appendix A.1.
The simple special case of $f_{X},f_{Y}>0$ in $\mathbb{R}$ illustrates the key ideas of Theorem 2.1. In this special case we have that:
@@
For simplicity of exposition, we continue with this special case where $f_{X},f_{Y}>0$ on all of $\mathbb{R}$, but note that adapting the results to the general case is a matter of technicality. Namely – $f_{X},f_{Y}>0$ on $S_{X},S_{Y}$ respectively. Combined with $S_{X}$ and $S_{Y}$ being open intervals, we conclude that the restricted CDFs ${F_{X}}|_{S_{X}},{F_{Y}}|_{S_{Y}}$ are continuous and strictly monotonically increasing on $S_{X},S_{Y}$, which is the main observation needed in order to generalize.
@@
4. Guiding topological properties – There are useful properties that are invariant under homeomorphisms (”topological properties”) that can guide us in designing better generative models. For example, one property is that homeomorphic manifolds must have the same dimension. The use of this property is demonstrated under Example: Swiss-Roll Parameter, where it will lead us to the conclusion that we are better off generating swiss-roll samples from a $1$-dimensional standard normal distribution, rather than from a $2$-dimensional one. Another example for a useful property is in Appendix A.3.
Figure 3: (A) Test results for a GTN $\hat{h}$ trained to map from $Y\sim\mathcal{N}(0,1)$ to the swiss-roll parameter. The color indicates which point in the normal sample was mapped to which point in the swiss-roll. (B) Test results for a GTN $\hat{h}$ trained to map from $Y\sim\mathcal{N}(0,I)$ to $X\sim U\big{(}(0,1)\times(0,1)\big{)}$. The color is based on the normal sample, i.e. for each $y$ in the normal sample, $\hat{h}(y)$ has the same color as $y$ so that the figure on the right shows how the normal sample was stretched to a uniform distribution (which is in line with Equation 4).
@@
Let $D_{X}:=\{x_{1},\dots,x_{n}\}$ and $D_{Y}:=\{y_{1},\dots y_{n}\}$ be $n$ observed values of $X$ and $Y$, correspondingly. Labels are obtained easily by sorting $D_{X}$ and $D_{Y}$ to obtain $D_{X}^{\text{sorted}}$ and $D_{Y}^{\text{sorted}}$ and labeling each $y\in D_{Y}^{\text{sorted}}$ with the corresponding $x\in D_{X}^{\text{sorted}}$ (i.e. assuming $D_{X},D_{Y}$ are sorted, $y_{i}$ is assigned $x_{i}$). This is illustrated in Figure 2.
@@
Example: Swiss-Roll Parameter.
In this example, we will see that samples from the swiss-roll can be easily generated with visibly near-perfect accuracy (Figure 3 (A)) using this method. This would not be as easy using a $2$-dimensional normal distribution (the approach taken by previous methods such as (Sohl-Dickstein et al. 2015a)), as we will now explain.
The swiss-roll is a $1$-dimensional manifold in a 2-dimensional space $\mathbb{R}^{2}$ (generated using a single, 1-dimensional, continuous random variable $\theta$). Being a $1$-dimensional manifold, it cannot be homermorphic to the support of a $2$-dimensional normal distribution ($\mathbb{R}^{2}$) since homeomorphic manifolds must have the same dimension. Therefore, there would be no hope of learning a homeomorphism that maps a 2-dimensional Gaussian random variable to the swiss-roll. However, the swiss roll is homeomorphic to $\mathbb{R}$ – the support of a $1$-dimensional standard normal distribution. In fact, since $\theta$ is a random variable that satisfies the conditions in Theorem 2.1, it is homeomorphic to $\mathbb{R}$ via $h$, which we can learn using $\hat{h}$.
To train $\hat{h}$, we created a dataset of $n=50,000$ samples from $\theta$, denoted $D_{X}$. We sampled $n$ samples from a $1$-dimensional standard normal distribution to obtain $D_{Y}$. We labeled each $y_{i}\in D_{Y}$ with its $x_{y_{i}}\in D_{X}$ as defined in the previous section. We used a standard feedforward neural network as $\hat{h}$ ($3$ layers of width $5$ with ReLU activation and batch normalization) and MSE as the loss function. Figure 3 (A) shows the result of testing the trained model $\hat{h}$ on a set of new samples $y_{1},\dots,y_{k}$ drawn from $\mathcal{N}(0,1)$ (each point is obtained by predicting $\hat{\theta}_{i}:=\hat{h}(y_{i})$ and applying the formula for the swiss roll to $\hat{\theta}_{i}$).
For comparison, we refer the reader to Figure 1 (middle) in (Sohl-Dickstein et al. 2015a) where a trained diffusion model was applied to points sampled from a standard 2-dimensional normal distribution. The points are gradually and stochastically moved back towards the swiss-roll manifold, but the result contains points that fall outside of this manifold. With $\hat{h}$, the generated points lie entirely within the data manifold. In fact, this holds true early in the training process, as shown in Appendix Figure 6.
@@ 2.2 Higher Dimensions
To generalize to more than one dimension, we reduce to the $1$-dimensional case by considering lines that pass through the origin. Briefly, we take the random variables obtained by restricting $X$ and $Y$ to each line, and apply the homeomorphism defined for the $1$-dimensional case to these random variables. We begin with a formal setup that is very similar to the $1$-dimensional case.
@@
Let $X=(X_{1},\dots,X_{d}),Y=(Y_{1},\dots,Y_{d})$ be multivariate random-variables with continuous joint probability density functions $f_{X}$, $f_{Y}$, and with supports $S_{X}$, $S_{Y}$, each of which is a product of $d$ open intervals in $\mathbb{R}$ (see Appendix A.3. to accommodate for other supports). For example, $X$ could be uniformly distributed with support $S_{X}=(0,1)\times(0,1)$ and $Y$ could be normally distributed with support $S_{Y}=\mathbb{R}\times\mathbb{R}$.
For distributions that are symmetric about the origin (e.g. standard normal), we can define the following homeomorphism (we assume for simplicity that $S_{Y},S_{X}=\mathbb{R}^{d}$):
@@
where $h_{1}:(0,\infty)\to(0,\infty)$ is the $1$-dimensional homeomorphism applied to the random variables $||Y||,||X||$.
We prove that this is indeed a homeomorphism in Appendix A.5. Note that this generalizes the $1$-dimensional case since for $d=1$ we get $y\mapsto h_{1}(y)$. Intuitively, $h$ can be seen as shrinking $y$ to the unit vector in the same direction, and stretching(or shrinking it to reach the resulting $x$. This stretching and shrinking occurs on the same line through the origin, shared by both $x$ and $y$, where $h_{1}$ guarantees that the resulting $x$ has the same quantile as $y$ on this line (i.e. when measured with respect to the random variables obtained by restricting $Y,X$ to this line). Another way of thinking about this is provided in Appendix A.6.
@@
We now turn to our labeling approach. Let $D_{X}$ and $D_{Y}$ be observed values from $X$ and $Y$. Ideally, we would use the $1$-dimensional labeling scheme on each line that passes through the origin by using $h_{1}$ on both sides of the line. In practice, however, the probability that any given line contains points from either $D_{X}$ or $D_{Y}$ is negligent. We therefore approximate this labeling approach using cosine-similarity. Specifically, we match $y$ with the $x$ that is closest in terms of both cosine-similarity (so that both $y,x_{y}$ are as close as possible to being on the same line and in the same direction from the origin) and distance from the origin (to produce the effect of $h_{1}(||y||)$). This is formally described in Algorithm 1. To see why Algorithm 1 approximates $h$, consider a line with an infinite sample from both $X$ and $Y$ in both directions from the origin. The maximal cosine-similarity over all $x\in D_{X}$ and a given $y$ on this line is $1$. Since the algorithm returns as $x_{y}$ an $x$ with maximum cosine-similarity, then $x_{y}$ will indeed be on the line and in the same direction as $y$. Since there are infinitely many such candidates $x$ with cosine-similarity $1$, it remains to choose the particular $x$ whose distance from the origin best approximates $h_{1}(||y||)$. Since both $D_{X}$ and $D_{Y}$ are sorted in ascending order by distance from the origin, and since matched pairs $(x_{y},y)$ are always removed, the algorithm will indeed return the particular $x$ from this pool of candidates whose distance from the origin will have a similar empirical quantile to that of $||y||$. We further elaborate on Algorithm 1 in Appendix A.2.
Using the labeled samples from Algorithm 1, we train $\hat{h}$ – a feedforward neural network – using MSE as the loss function. We then use $\hat{h}$ to generate new $X$ instances just as in the $1$-dimensional case and as is formally described in Algorithm 2.
Note that the requirement of respecting the $1$-dimensional homeomorphisms on each line facilitates the desired continuous interpolations of the form described in the Method section. This advantage is later demonstrated in Figure 5.
@@
$\quad\quad\quad D_{Y}=\{y_{1},\dots,y_{n}\}\text{ sampled from }\mathcal{N}(0,I)$
@@
1:  $y\leftarrow\text{ sample from }\mathcal{N}(0,I)$
@@
Figure 3 (B) shows a sample generated using Algorithm 2 after applying the multivariate method to the multivariate uniform distribution $X\sim U(0,1)\times U(0,1)$. Specifically, we created $D_{X}$ by sampling $n=100,000$ points from $X$, and created $D_{Y}$ by sampling the same number of points from $\mathcal{N}(0,I)$. We then applied Algorithm 1 to $D_{X},D_{Y}$ to generate labeled data. We trained a standard feedforward neural network $\hat{h}$ ($5$ hidden layers of width $6$ with ReLU activation and batch normalization (Ioffe and Szegedy 2015)) and used it to generate new samples according to Algorithm 2. The model was trained until convergence on a separately generated validation set with $n=10,000$. The colors in both images in Figure 3 (B) are based on the distance of the points from the origin in the Gaussian sample (the left image in (B)), so that the right image reflects where each point in the Gaussian was predicted to ’move’ to by $\hat{h}$.
Figure 4: Randomly chosen samples generated by GTN during the training process on MNIST at four points of improvement on the validation set. Each epoch shows 100 images, each of which is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(0,I)$). The rightmost grid contains 100 images each of which is the decoded latent vector of a random real image.

### v2

@@ Abstract
Generative models have seen significant advancements in recent years, yet often remain challenging and costly to train and use. We introduce Generative Topological Networks (GTNs) – a new class of generative models that addresses these shortcomings. GTNs are trained deterministically using a simple supervised learning approach grounded in topology theory. GTNs are fast to train, and require only a single forward pass in a standard feedforward neural network to generate samples. We demonstrate the strengths of GTNs on several datasets, including MNIST, CelebA and the Hands and Palm Images dataset. Finally, the theory behind GTNs offers insights into how to train generative models for improved performance. Code and weights are available at: https://github.com/alonalj/GTN
@@ 1 Introduction
Figure 1: Samples generated by a GTN trained on CelebA $64\times 64$ with a latent dimension of $100$.
Deep generative models such as Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), Variational Autoencoders (VAEs) (Kingma and Welling 2013), Energy-Based Models (EBMs) (LeCun et al. 2006; Ngiam et al. 2011), normalizing flows (NFs) (Rezende and Mohamed 2015), and diffusion models (Sohl-Dickstein et al. 2015b), have demonstrated remarkable capabilities for generating samples based on training data distributions (Kang et al. 2023; Ho, Jain, and Abbeel 2020; Ramesh et al. 2021; Ho et al. 2019; Kingma and Dhariwal 2018; Kingma et al. 2016; Reed et al. 2017; Van Den Oord, Vinyals et al. 2017; Ramesh et al. 2021; Onken et al. 2021).
@@
However, diffusion models have several downsides. Most prominent is their inferior time-efficiency during training and sampling when compared to VAEs and GANs. This is because they employ a gradual and stochastic transition from a Gaussian distribution to the data distribution, performed over multiple time-steps known as the reverse diffusion process. The complexity of this process also makes it challenging to produce interpolations – desirable in the context of continuous data like audio and video. Although recent advancements have made significant steps towards enhancing diffusion models, the complex diffusion process is still a key challenging component (Manduchi et al. 2024).
@@
From a practical perspective, GTNs are extremely simple and efficient, requiring a single, vanilla, feedforward architecture trained using standard supervised learning. This means that GTNs avoid the complexities of introducing stochasticity during training which is present in VAE and diffusion; circumvent the intricacies of training more complex architectures such as those employed by diffusion and GANs; are not a likelihood-based approach, as are most generative models; and do not pose constraints on the structure of the neural network, as do NFs. GTNs also operate on latent representations of the data, providing even more enhanced training and sampling efficiency. These advantages manifest in our experiments – realistic samples are obtained at early training epochs, and generating a single image takes around $4$ milliseconds on a standard PC and $0.003$ milliseconds on a T4 GPU.
@@
The remainder of the paper is structured as follows: In the Method section we first develop the theory behind GTNs for the $1$-dimensional (1D) case and use it to generate samples from the swiss-roll. We then extend this theory to higher dimensions and use it to generate samples from the multivariate uniform distribution. In the Experiments section we apply GTNs to real datasets, also including demonstrations of continuous interpolation. We follow with Related Work section and conclude with the Discussion section.
@@ 2 Method
We first develop the method in the simple case of a $1$D space, and proceed to generalize to higher dimensions.
@@
In this section we will define the aforementioned function $h$ that transforms one distribution into another. We will also prove that $h$, under fairly general conditions, possesses certain properties that are desirable in the context of generative models by proving that it is a homeomorphism (Theorem 2.1). For example, we will see that $h$ is bijective, so that each sample $y$ is mapped precisely to one sample $x_{y}=h(y)$, generating different samples for different $y$, and guaranteeing that each sample $x$ has a sample $y$ that generates it. We will also discuss other useful consequences of $h$ being a homeomorphism. We begin by defining the term homeomorphism in our context and proceed to defining $h$ in Theorem 2.1.
Definition 2.1 (Homeomorphism for $\mathbb{R}^{n}$). Let $S,T$ be two subsets of $\mathbb{R}^{n}$. A function $h:S\to T$ is a homeomorphism if: (1) $h$ is continuous; (2) $h$ is bijective; (3) $h^{-1}$ is continuous. When such an $h$ exists, then $S$ and $T$ are called homeomorphic.
Theorem 2.1 Let $X$ and $Y$ be random variables that have continuous probability density functions (pdfs) $f_{X}$, $f_{Y}$ and supports $S_{X}$, $S_{Y}$ that are open intervals in $\mathbb{R}$. Denote the corresponding cumulative distribution functions (CDFs) as $F_{X}$ and $F_{Y}$. Define:
@@
Note that the requirement that $S_{X},S_{Y}$ are open intervals can be adapted to a union of open intervals – see Appendix C. The proof of Theorem 2.1 is in Appendix A.
The simple special case of $f_{X},f_{Y}>0$ in $\mathbb{R}$ illustrates the key ideas of Theorem 2.1. In this case we have that:
@@
For simplicity of exposition, we continue with this special case but note that adapting the results to the general case is a matter of technicality. Namely – $f_{X},f_{Y}>0$ on $S_{X},S_{Y}$ respectively. Combined with $S_{X}$ and $S_{Y}$ being open intervals, we conclude that the restricted CDFs ${F_{X}}|_{S_{X}},{F_{Y}}|_{S_{Y}}$ are continuous and strictly monotonically increasing on $S_{X},S_{Y}$, which is the main observation needed in order to generalize.
@@
4. Guiding topological properties – There are useful properties that are invariant under homeomorphisms (”topological properties”) that can guide us in designing better generative models. For example, one property is that homeomorphic manifolds must have the same dimension. The use of this property is demonstrated in the upcoming swiss-roll example, where it will lead us to the conclusion that we are better off generating swiss-roll samples from a $1$D standard normal distribution, rather than from a $2$D one. Another example for a useful property is in Appendix C.
Figure 3: (A) Test results for a GTN $\hat{h}$ trained to map from $Y\sim\mathcal{N}(0,1)$ to the swiss-roll parameter. The color indicates which point in the normal sample was mapped to which point in the swiss-roll. (B) Test results for a GTN $\hat{h}$ trained to map from $Y\sim\mathcal{N}(\textbf{0},\textbf{I})$ to $X\sim U\big{(}(0,1)\times(0,1)\big{)}$. The color is based on the normal sample (left): for each $y$ in the normal sample, $\hat{h}(y)$ has the same color as $y$ so that the figure on the right shows how the normal sample was stretched to a uniform distribution.
@@
Let $D_{X}:=\{x_{1},\dots,x_{n}\}$ and $D_{Y}:=\{y_{1},\dots y_{n}\}$ be $n$ observed values of $X$ and $Y$, respectively. Labels are obtained easily by sorting $D_{X}$ and $D_{Y}$ to obtain $D_{X}^{\text{sorted}}$ and $D_{Y}^{\text{sorted}}$ and labeling each $y\in D_{Y}^{\text{sorted}}$ with the corresponding $x\in D_{X}^{\text{sorted}}$ (i.e. assuming $D_{X},D_{Y}$ are sorted, $y_{i}$ is assigned $x_{i}$). This is illustrated in Figure 2.
@@
Example: Swiss-Roll.
In this example, we will see that samples from the swiss-roll can be easily generated from a $1$D normal distribution with visibly near-perfect accuracy (Figure 3 (A)) using a GTN. This would not be as easy using a $2$D normal distribution (the approach taken by previous methods such as (Sohl-Dickstein et al. 2015b)), as we will now explain.
The swiss-roll is a $1$D manifold in a $2$D space $\mathbb{R}^{2}$ (generated using a single, 1-dimensional, continuous random variable $\theta$). Being a $1$D manifold, it cannot be homermorphic to the support of a $2$D normal distribution ($\mathbb{R}^{2}$) since homeomorphic manifolds must have the same dimension. Therefore, there would be no hope of learning a homeomorphism that maps a $2$D Gaussian random variable to the swiss-roll. However, the swiss roll is homeomorphic to $\mathbb{R}$ – the support of a $1$D standard normal distribution. In fact, since $\theta$ is a random variable that satisfies the conditions in Theorem 2.1, it is homeomorphic to $\mathbb{R}$ via $h$, which we can learn using $\hat{h}$. The fact that there is no hope of learning a homeomorphism in a higher-dimensional space than the intrinsic dimension is also important for NFs, which use a diffeomorphism (a type of homeomorphism) as a generator, and it may also help explain why latent diffusion models have shown improvements over pixel-space diffusion models (Rombach et al. 2022).
To train $\hat{h}$, we created a dataset of $n=50,000$ samples from $\theta$, denoted $D_{X}$. We sampled $n$ samples from a $1$D standard normal distribution to obtain $D_{Y}$. We labeled each $y_{i}\in D_{Y}$ with its $x_{y_{i}}\in D_{X}$ as defined in the previous section. We used a standard feedforward neural network as $\hat{h}$ ($4$ layers of width $6$; details in Appendix Table 1). Figure 3 (A) shows the result of testing the trained model $\hat{h}$ on a set of new samples $y_{1},\dots,y_{k}$ drawn from $\mathcal{N}(0,1)$ (each point is obtained by predicting $\hat{\theta}_{i}:=\hat{h}(y_{i})$ and applying the formula for the swiss roll to $\hat{\theta}_{i}$).
For comparison, we refer the reader to Figure 1 (middle) in (Sohl-Dickstein et al. 2015b) where a trained diffusion model was applied to points sampled from a standard $2$D normal distribution. The points are gradually and stochastically moved back towards the swiss-roll manifold, but the result contains points that fall outside of this manifold. With $\hat{h}$, the generated points lie entirely within the data manifold. In fact, this holds true early in the training process, as shown in Appendix Figure 8.
@@ 2.2 Higher Dimensions
To generalize to more than one dimension, we reduce to the $1$D case by considering lines that pass through the origin. Briefly, we take the random variables obtained by restricting $X$ and $Y$ to each line, and apply the homeomorphism defined for the $1$D case to these random variables. We begin with a formal setup that is very similar to the $1$D case.
@@
Let $X=(X_{1},\dots,X_{d}),Y=(Y_{1},\dots,Y_{d})$ be multivariate random-variables with continuous joint probability density functions $f_{X}$, $f_{Y}$, and with supports $S_{X}$, $S_{Y}$, each of which is a product of $d$ open intervals in $\mathbb{R}$ (see Appendix C to accommodate for other supports). For example, $X$ could be uniformly distributed with support $S_{X}=(0,1)\times(0,1)$ and $Y$ could be normally distributed with support $S_{Y}=\mathbb{R}\times\mathbb{R}$.
Consider first rotation invariant distributions (e.g., the standard multivariate normal). That is, assume that for every rotation matrix $R$ and every $x\in\mathbb{R}^{d}$, $f_{X}(Rx)=f_{X}(x)$, and similarly for $f_{Y}$. For simplicity, also assume that $S_{Y},S_{X}=\mathbb{R}^{d}$. We can now define the following homeomorphism:
@@
where $h_{1}:(0,\infty)\to(0,\infty)$ is the $1$D homeomorphism applied to the random variables $||Y||,||X||$.
We prove that this is indeed a homeomorphism in Appendix D. Note that this generalizes the $1$D case since for $d=1$ we get $y\mapsto h_{1}(y)$. Intuitively, $h$ can be seen as shrinking or stretching $y$ to the unit vector in $y$’s direction ($y/||y||)$, and then shrinking or stretching it to reach the $x_{y}:=h(y)$ that has the same ranked distance as $y$ on the line (by multiplying it by $h_{1}(||y||)$). More precisely, $h_{1}$ produces $x_{y}$’s distance from the origin so that it has the same quantile as $y$’s distance from the origin (when measured with respect to the random variables obtained by restricting $Y,X$ to the line segment from the origin in $y$’s direction). Another way of thinking about this is provided in Appendix E.
For more complicated distributions, it may be difficult to explicitly define a similar $h$ since $h_{1}$ would depend on the line. Nevertheless, the simpler case above can guide us on how to train a neural network to learn one (as accomplished in Figure 3 (B)).
@@
Let $D_{X}$ and $D_{Y}$ be observed values from $X$ and $Y$. Ideally, we would use the $1$D labeling scheme on each line that passes through the origin by using $h_{1}$ on both sides of the line. In practice, however, the probability that any given line contains points from either $D_{X}$ or $D_{Y}$ is negligible. We therefore approximate this labeling approach using cosine-similarity. Specifically, we match $y$ with the $x$ that is closest in terms of both cosine-similarity (so that both $y,x_{y}$ are as close as possible to being on the same line and in the same direction) and distance from the origin (to approximate $h_{1}(||y||)$). This is formally described in Algorithm 1 and illustrated in Figure 4. To see why Algorithm 1 approximates $h$, consider a line with an infinite sample from $X$ and $Y$ in both directions from the origin (see Figure 4). The maximal cosine-similarity over all $x\in D_{X}$ and a $y$ on this line is $1$, so the algorithm returns an $x_{y}$ that is on this line and in $y$’s direction. Because $D_{X},D_{Y}$ are sorted by distance from the origin, ties in cosine-similarity are broken by distance from the origin, so that the first $y$ is matched with the first $x$, the second $y$ with the second $x$ etc. as in the $1$D case, i.e. $||x_{y}||$ approximates $h_{1}(||y||)$. Appendix B gives further intuition.
Using the labeled samples from Algorithm 1, we train $\hat{h}$ – a feedforward neural network – using MSE as the loss function. We then use $\hat{h}$ to generate new $X$ instances just as in the $1$D case, formally described in Algorithm 2.
@@
$\quad\quad\quad D_{Y}=\{y_{1},\dots,y_{n}\}\text{ sampled from }\mathcal{N}(\textbf{0},\textbf{I})$
@@
1:  $y\leftarrow\text{ sample from }\mathcal{N}(\textbf{0},\textbf{I})$
@@
Figure 4: Illustration of Algorithm 1. Numbers reflect the order of matching $y$ (circles) with $x_{y}$ (stars).
@@
Figure 3 (B) shows a sample generated using Algorithm 2 after applying the method to the multivariate uniform distribution $X\sim U(0,1)\times U(0,1)$. Specifically, we created $D_{X}$ by sampling $n=100,000$ points from $X$, and created $D_{Y}$ by sampling the same number of points from $\mathcal{N}(\textbf{0},\textbf{I})$. We then applied Algorithm 1 to $D_{X},D_{Y}$ to generate labeled data. We trained a standard feedforward neural network $\hat{h}$ (6 layers, width 6; details in Appendix Table 1) and used it to generate new samples according to Algorithm 2. The model was trained until convergence on a separately generated validation set with $n=10,000$. The colors in both images in Figure 3 (B) are based on the distance of the points from the origin in the Gaussian sample (the left image in (B)), so that the image on the right reflects where each point in the Gaussian sample was predicted to ’move’ to by $\hat{h}$.
Figure 5: Randomly chosen samples generated by GTN during the training process on MNIST at four points of improvement on the validation set. Each epoch shows 100 images, each of which is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). The rightmost grid contains 100 images each of which is the decoded latent vector of a random real image.